3(x/2)+5=19
x=???
AND
24x+2-5x+42=12
x= ??
thx !

Answer by TimothyLamb(4379)

3(x/2)+5=19
24x+2-5x+42=12
---
(3/2)x + 5 = 19
24x + 2 - 5x + 42 = 12
---
(3/2)x = 14
19x = -32
---
x = 14(2/3)
x = -32/19
---
answer:
x = 9.333
x = -1.684
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php